# Advanced User Credit Risk Prediction Model using LightGBM, XGBoost and Tabnet with SMOTEENN

### v1

@@ Abstract
Bank credit risk is a significant challenge in modern financial transactions, and the ability to identify qualified credit card holders among a large number of applicants is crucial for the profitability of a bank’sbank’s credit card business. In the past, screening applicants’applicants’ conditions often required a significant amount of manual labor, which was time-consuming and labor-intensive. Although the accuracy and reliability of previously used ML models have been continuously improving, the pursuit of more reliable and powerful AI intelligent models is undoubtedly the unremitting pursuit by major banks in the financial industry. In this study, we used a dataset of over 40,000 records provided by a commercial bank as the research object. We compared various dimensionality reduction techniques such as PCA and T-SNE for preprocessing high-dimensional datasets and performed in-depth adaptation and tuning of distributed models such as LightGBM and XGBoost, as well as deep models like Tabnet. After a series of research and processing, we obtained excellent research results by combining SMOTEENN with these techniques. The experiments demonstrated that LightGBM combined with PCA and SMOTEENN techniques can assist banks in accurately predicting potential high-quality customers, showing relatively outstanding performance compared to other models.
@@ I Introduction
Bank operations face many risks, among which credit risk is the primary concern. Credit risk caused by the possibility of loss to creditors due to the debtor’sdebtor’s failure to fulfill obligations as per the contract or changes in credit quality[46, 52]. Credit risk can cause direct or indirect economic losses to banks, increase management costs, affect the utilization of bank funds, and seriously hinder the bank’sbank’s lending operations. Based on the above reasons, establishing a credit risk prediction model can help customers predict whether they will default based on their data information, which can assist banks in controlling risks, reducing losses[45, 13], and increasing potential revenue.
In this paper, we compare various classical models used in major banks, such as the prediction model based on Random Forest and Logistic Regression used by Milad Malkipirbazar on social credit platforms, which achieved a risk prediction accuracy of 88% for Random Forest and 49% for Logistic Regression [37]. Later models, such as Neural Networks[18], also achieved an accuracy of over 80% for the fraud data they used [12]. Ke, G., and Meng, in their paper ”LightGBM: A Highly Efficient Gradient Boosting Decision Tree. Advances in Neural Information Processing Systems,” mentioned the theoretical application of using LightGBM for high-performance classification and processing [20]. Based on the principles they proposed, we combined PCA dimensionality reduction and SMOTEENN composite sampling methods and achieved very promising success through a series of data processing steps[34].
This study focuses on building more powerful models to predict User Credit Risk. We conducted multiple model training sessions based on the original data[17]. First, we tuned a series of algorithms, including LightGBM, XGBoost, CatBoost, Tabnet, and Neural Network, on the original data. After the initial screening, we combined PCA dimensionality reduction techniques with these models to optimize their performance[49]. On this basis, we incorporated the advanced SMOTEENN technique, which greatly improved the performance of XGBoost, LightBoost, CatBoost, and Tabnet, with LightGBM showing particularly outstanding performance.
@@ II Related Work
Guolin Ke conducted a study on using machine learning for credit risk prediction, proposing the necessity for financial institutions to utilize artificial intelligence and machine learning to assess credit risk and analyze vast amounts of information. The author raised research questions regarding the algorithms, evaluation metrics, results, datasets, variables, and related limitations used in credit risk prediction, and found 52 relevant studies in the microcredit industry through searching well-known databases [20, 23].
@@
Abdou, H. A., and Pointon, J.’s article summarizes the current state of research in the field of credit scoring. It provides a comprehensive introduction and discussion of commonly used statistical techniques, evaluation criteria, and data processing methods. Their article serves as an excellent reference for researchers in the field of credit scoring and offers valuable guidance for financial institutions in developing and optimizing credit scoring models [2, 51].
@@
Li et al. [27] proposed using self-supervised local features and quantization to improve the tracking stability and real-time performance of feature-based VSLAM under illumination and viewpoint changes. Their approach to feature quantization has inspired us to explore similar techniques for optimizing our data processing pipeline, leading to more efficient and robust data analysis.we also find[17] that visual prompting can help models find better sparse networks and propose a novel data-model co-design sparsification method.
@@ III-A Data Analysis
Our approach involved a detailed assessment of the IV values associated with various features. IV is a metric used to quantify the predictive power of a variable in relation to the target variable. It helps identify the most significant features of the model. Features with higher IV values are considered more predictive and, thus, more important for the model’s performance. The work [16, 8] rethought and designed a new paradigm for data preprocessing that also significantly improves the model efficiency with neglectable performance degradation.
By calculating the IV for each feature, we were able to prioritize them based on their contributions to the target variable. We find that generalizable correlations are easier to be imitated by other learners and propose a novel regularization method to enhance the generalization of CNNs-based, RNNs-based, and Transformer-based models[10, 35]. This prioritization allowed us to focus on the most relevant features during the training process, thereby improving both the efficiency of the training phase and the accuracy of the resulting model as Fig below.
@@ III-B Random UnderSamping
In our original experimental dataset, the number of ”not approve” versus ”approve” instances was 45,318 to 667. Given this extreme imbalance, we applied Random UnderSampling to balance the sample sizes of each class. By randomly under-sampling the majority class, we were able to equalize the number of samples across classes, enabling the model to learn the features of each class more fairly[50, 7]. This reduction in dataset size also accelerated the training process and conserved computational resources.
Random UnderSampling involves multiple iterations of sampling and training. By repeatedly creating different subsets of the data and training the model on these subsets, we obtained a set of models based on various data samples. We then employed ensemble learning techniques to combine these models, thereby enhancing overall performance.Inspired by the work of Liu et al. [32], we have adopted a similar approach to efficiently analyze and process our data. Their method has significantly optimized our processing pipeline, laying a solid foundation for subsequent program processes.
@@ III-C Dimensionality Reduction using Principal Component Analysis
We utilized Principal Component Analysis (PCA), our dimensionality reduction technique, in this study. PCA is a widely used statistical technology that changes high-dimensional data into a lower-dimensional space while retaining the essential information [19]. The primary goal of PCA is to identify the principal components, which are linear combinations of the original features that capture the maximum variance in the data [1].
@@
Furthermore, PCA enables data visualization and exploratory analysis by projecting high-dimensional data into a lower-dimensional space, typically two or three dimensions. Visualizing the PCA results allows for a more intuitive understanding of the data’s intrinsic structure, clustering patterns, and relationships between different categories [42]. This facilitates data interpretation and aids in gaining insights from complex datasets.
@@ III-D Synthetic Minority Over-sampling Technique and Edited Nearest Neighbors
Combination of Over-sampling and Under-sampling: SMOTEENN combines SMOTE over-sampling and ENN under-sampling to balance the dataset by increasing the number of minority class samples and decreasing the number of majority class samples[22].
@@ III-G TabnNet
TabNet, introduced by Google, is a neural network model designed for tabular data. It uses a sequential attention mechanism for feature selection and an encoder-decoder framework for self-supervised learning. Compared to traditional tree-based models like Random Forest and XGBoost, TabNet offers advantages in performance, interpretability, and generalization. The core idea involves sequentially applying an Attentive Transformer for feature selection and a Feature Transformer for converting selected features into useful representations. This process mimics the decision-making of trees and is repeated for multiple steps, controlled by a hyperparameter. TabNet also incorporates sparsity and feature reuse to enhance performance and prevent overfitting, using prior scales to control feature selection frequency and entropy-based loss functions to make attention masks sparser. The model provides both local and global interpretability, allowing visualization of feature importance and overall feature contributions. Through these mechanisms, TabNet constructs a complex nonlinear transformation in the feature space, enhancing expressive power while maintaining interpretability.
@@ IV-A Evaluation Metrics
Precision will measure the proportion of actual positive instances among those predicted as positive by the model [39, 29]. It reflects the model’s accuracy in correctly identifying positive instances, calculated as:
@@
On the other hand, Recall represents the proportion of actual positive instances correctly identified by the model [15, 21]. It captures the model’s ability to identify all relevant instances, defined as:
@@
The F1 Score is generated by the harmonic mean of Precision and Recall, which provides a balanced measure that considers the accuracy and the coverage of the model [31, 40]. It is calculated as:
@@ IV-B1 Origin Data
As evident from Table I, the performance of the models on the raw data was not satisfactory, with F1 scores ranging from 0.4652 to 0.6388. LightGBM achieved the highest F1 score of 0.6388, while TabNet had the lowest at 0.4652. The AUC-ROC values also indicated poor discriminatory power[38, 25], with the highest being 0.6593 for Random Forest and the lowest being 0.5207 for NGBoost. These results suggest that the raw data alone may not be sufficient for building a highly accurate predictive model, and further preprocessing and feature engineering may be necessary.
@@ IV-B2 Model Performance with PCA
In the preliminary processing, we observed that the accuracy of the models was generally low. To improve the model performance, we employed Principal Component Analysis for dimensionality reduction of the data[11, 48]. After applying PCA, the models exhibited exceptional performance, and the results were remarkable, as shown in Table II.
@@ IV-B3 Further Optimization with SMOTEENN
By comparing the results across different scenarios, we can observe that LightGBM consistently outperformed other models[9], even without PCA or SMOTEENN. In all environments, LightGBM achieved the highest performance metrics. When combined with both SMOTEENN and PCA techniques, LightGBM demonstrated an exceptionally high accuracy, with an F1 score and AUC-ROC of 0.9989 and 0.9999, respectively. These experiment results are highly encouraging and highlight the superiority of LightGBM in handling the given dataset and task.
Conclusion
The research work on User Risk Prediction has always been a top priority in the financial domain[29]. Through a series of experiments, we discovered that LightGBM, Random Forest, and CatBoost perform exceptionally well in predicting credit risk. Among these models, LightGBM demonstrated outstanding performance, showcasing its vast potential for applications in this field.
Our study investigated different data processing methods and their effects on model performance[47]. We discovered that combining Principal Component Analysis (PCA) for dimensionality reduction with SMOTEENN for addressing imbalanced datasets significantly enhanced the models’ predictive capabilities. By implementing these techniques, we achieved notably high accuracy in User Risk Prediction.
@@ Conclusion
Our findings have significant implications for financial institutions and businesses involved in credit risk assessment[43]. By leveraging machine learning models like LightGBM, Random Forest, and CatBoost, together with appropriate data processing techniques, organizations can enhance their risk management strategies and make more informed decisions[30].
@@
In conclusion, our research demonstrates the immense potential of machine learning techniques, particularly LightGBM, in User Risk Prediction[26, 44]. The combination of advanced data processing methods and powerful predictive models opens up a wide range of applications in the financial domain. By embracing these approaches, organizations can significantly improve their risk assessment capabilities and make data-driven decisions to mitigate potential losses and optimize their operations.

### v2

@@ Abstract
Bank credit risk is a significant challenge in modern financial transactions, and the ability to identify qualified credit card holders among a large number of applicants is crucial for the profitability of a bank’s credit card business. In the past, screening applicants’ conditions often required a significant amount of manual labor, which was time-consuming and labor-intensive. Although the accuracy and reliability of previously used ML models have been continuously improving, the pursuit of more reliable and powerful AI intelligent models is undoubtedly the unremitting pursuit by major banks in the financial industry. In this study, we used a dataset of over 40,000 records provided by a commercial bank as the research object. We compared various dimension reduction techniques such as PCA and T-SNE for processing high-dimensional datasets and performed in-depth adaptation and tuning of distributed models such as LightGBM and XGBoost, as well as deep models like Tabnet. After a series of research and processing, we obtained excellent research results by combining SMOTEENN with these techniques. The experiments demonstrated that LightGBM combined with PCA and SMOTEENN techniques can assist banks in accurately predicting potential high-quality customers, showing relatively outstanding performance compared to other models.
@@ I Introduction
Bank operations face many risks, among which credit risk is the primary concern. Credit risk caused by the possibility of loss to creditors due to the debtor’s failure to fulfill obligations as per the contract or changes in credit quality. Credit risk can cause direct or indirect economic losses to banks, increase management costs, affect the utilization of bank funds, and seriously hinder the bank’s lending operations. Based on the above reasons, establishing a credit risk prediction model can help customers predict whether they will default based on their data information, which can assist banks in controlling risks, reducing losses, and increasing potential revenue.
In this paper, we compare various classical models used in major banks, such as the prediction model based on Random Forest and Logistic Regression used by Milad Malkipirbazar on social credit platforms, which achieved a risk prediction accuracy of 88% for Random Forest and 49% for Logistic Regression [16]. Later models, such as Neural Networks[9], also achieved an accuracy of over 80% for the fraud data they used [7]. Ke, G., and Meng, in their paper ”LightGBM: A Highly Efficient Gradient Boosting Decision Tree. Advances in Neural Information Processing Systems,” mentioned the theoretical application of using LightGBM for high-performance classification and processing. Based on the principles they proposed, we combined PCA dimensionality reduction and SMOTEENN composite sampling methods and achieved very promising success through a series of data processing steps.
This study focuses on building more powerful models to predict User Credit Risk. We conducted multiple model training sessions based on the original data. First, we tuned a series of algorithms, including LightGBM, XGBoost, CatBoost, Tabnet, and Neural Network, on the original data. After the initial screening, we combined PCA dimensionality reduction techniques with these models to optimize their performance. On this basis, we incorporated the advanced SMOTEENN technique, which greatly improved the performance of XGBoost, LightBoost, CatBoost, and Tabnet, with LightGBM showing particularly outstanding performance.
@@ II Related Work
Guolin Ke conducted a study on using machine learning for credit risk prediction, proposing the necessity for financial institutions to utilize artificial intelligence and machine learning to assess credit risk and analyze vast amounts of information. The author raised research questions regarding the algorithms, evaluation metrics, results, datasets, variables, and related limitations used in credit risk prediction, and found 52 relevant studies in the microcredit industry through searching well-known databases [ke2017lightgbmg].
@@
Abdou, H. A., and Pointon, J.’s article summarizes the current state of research in the field of credit scoring. It provides a comprehensive introduction and discussion of commonly used statistical techniques, evaluation criteria, and data processing methods. Their article serves as an excellent reference for researchers in the field of credit scoring and offers valuable guidance for financial institutions in developing and optimizing credit scoring models [2].
@@
Li et al. [11] proposed using self-supervised local features and quantization to improve the tracking stability and real-time performance of feature-based VSLAM under illumination and viewpoint changes. Their approach to feature quantization has inspired us to explore similar techniques for optimizing our data processing pipeline, leading to more efficient and robust data analysis.we also find that visual prompting can help models find better sparse networks and propose a novel data-model co-design sparsification method.
@@ III-A Data Analysis
Our approach involved a detailed assessment of the IV values associated with various features. IV is a metric used to quantify the predictive power of a variable in relation to the target variable. It helps identify the most significant features of the model. Features with higher IV values are considered more predictive and, thus, more important for the model’s performance. The work rethought and designed a new paradigm for data preprocessing that also significantly improves the model efficiency with neglectable performance degradation.
By calculating the IV for each feature, we were able to prioritize them based on their contributions to the target variable.The corresponding features are listed with their IV in below Fig 1. We find that generalizable correlations are easier to be imitated by other learners and propose a novel regularization method to enhance the generalization of CNNs-based, RNNs-based, and Transformer-based models. This prioritization allowed us to focus on the most relevant features during the training process, thereby improving both the efficiency of the training phase and the accuracy of the resulting model as Fig below.
@@ III-B Random Undersamping
In our original experimental dataset, the number of ”not approve” versus ”approve” instances was 45,318 to 667. Given this extreme imbalance, we applied Random Undersampling to balance the sample sizes of each class. By randomly under-sampling the majority class, we were able to equalize the number of samples across classes, enabling the model to learn the features of each class more fairly. This reduction in dataset size also accelerated the training process and conserved computational resources.
Random UnderSampling involves multiple iterations of sampling and training. By repeatedly creating different subsets of the data and training the model on these subsets, we obtained a set of models based on various data samples. We then employed ensemble learning techniques to combine these models, thereby enhancing overall performance.Inspired by the work of Liu et al [13], we have adopted a similar approach to efficiently analyze and process our data. Their method has significantly optimized our processing pipeline, laying a solid foundation for subsequent program processes.
@@ III-C Dimensionality Reduction using Principal Component Analysis
We utilized Principal Component Analysis (PCA), our dimensionality reduction technique, in this study. PCA is a widely used statistical technology that changes high-dimensional data into a lower-dimensional space while retaining the essential information . The primary goal of PCA is to identify the principal components, which are linear combinations of the original features that capture the maximum variance in the data [1].
@@
Furthermore, PCA enables data visualization and exploratory analysis by projecting high-dimensional data into a lower-dimensional space, typically two or three dimensions. Visualizing the PCA results allows for a more intuitive understanding of the data’s intrinsic structure, clustering patterns, and relationships between different categories . This facilitates data interpretation and aids in gaining insights from complex datasets.
@@ III-D Synthetic Minority Over-sampling Technique and Edited Nearest Neighbors
Combination of Over-sampling and Under-sampling: SMOTEENN combines SMOTE over-sampling and ENN under-sampling to balance the dataset by increasing the number of minority class samples and decreasing the number of majority class samples.
@@ III-G TabnNet
TabNet, introduced by Google, is a neural network model designed for tabular data. It uses a sequential attention mechanism for feature selection and an encoder-decoder framework for self-supervised learning. Compared to traditional tree-based models like Random Forest and XGBoost, TabNet offers advantages in performance, interpret-ability, and generalization. The core idea involves sequentially applying an Attentive Transformer for feature selection and a Feature Transformer for converting selected features into useful representations. This process mimics the decision-making of trees and is repeated for multiple steps, controlled by a hyper-parameter. TabNet also incorporates sparsity and feature reuse to enhance performance and prevent overfitting, using prior scales to control feature selection frequency and entropy-based loss functions to make attention masks sparser. The model provides both local and global interpret-ability, allowing visualization of feature importance and overall feature contributions. Through these mechanisms, TabNet constructs a complex nonlinear transformation in the feature space, enhancing expressive power while maintaining interpret-ability.
@@ IV-A Evaluation Metrics
Precision will measure the proportion of actual positive instances among those predicted as positive by the model . It reflects the model’s accuracy in correctly identifying positive instances, calculated as:
@@
On the other hand, Recall represents the proportion of actual positive instances correctly identified by the model . It captures the model’s ability to identify all relevant instances, defined as:
@@
The F1 Score is generated by the harmonic mean of Precision and Recall, which provides a balanced measure that considers the accuracy and the coverage of the model . It is calculated as:
@@ IV-B1 Origin Data
As evident from Table I, the performance of the models on the raw data was not satisfactory, with F1 scores ranging from 0.4652 to 0.6388. LightGBM achieved the highest F1 score of 0.6388, while TabNet had the lowest at 0.4652. The AUC-ROC values also indicated poor discriminatory power, with the highest being 0.6593 for Random Forest and the lowest being 0.5207 for NGBoost. These results suggest that the raw data alone may not be sufficient for building a highly accurate predictive model, and further preprocessing and feature engineering may be necessary.
@@ IV-B2 Model Performance with PCA
In the preliminary processing, we observed that the accuracy of the models was generally low. To improve the model performance, we employed Principal Component Analysis for dimensionality reduction of the data. After applying PCA, the models exhibited exceptional performance, and the results were remarkable, as shown in Table II.
@@ IV-B3 Further Optimization with SMOTEENN
By comparing the results across different scenarios, we can observe that LightGBM consistently outperformed other models, even without PCA or SMOTEENN. In all environments, LightGBM achieved the highest performance metrics. When combined with both SMOTEENN and PCA techniques, LightGBM demonstrated an exceptionally high accuracy, with an F1 score and AUC-ROC of 0.9989 and 0.9999, respectively. These experiment results are highly encouraging and highlight the superiority of LightGBM in handling the given dataset and task.
V Conclusion
The research work on User Risk Prediction has always been a top priority in the financial domain. Through a series of experiments, we discovered that LightGBM, Random Forest, and CatBoost perform exceptionally well in predicting credit risk. Among these models, LightGBM demonstrated outstanding performance, showcasing its vast potential for applications in this field.
Our study investigated different data processing methods and their effects on model performance. We discovered that combining Principal Component Analysis (PCA) for dimensionality reduction with SMOTEENN for addressing imbalanced datasets significantly enhanced the models’ predictive capabilities. By implementing these techniques, we achieved notably high accuracy in User Risk Prediction.
@@ V Conclusion
Our findings have significant implications for financial institutions and businesses involved in credit risk assessment. By leveraging machine learning models like LightGBM, Random Forest, and CatBoost, together with appropriate data processing techniques, organizations can enhance their risk management strategies and make more informed decisions.
@@
In conclusion, our research demonstrates the immense potential of machine learning techniques, particularly LightGBM, in User Risk Prediction. The combination of advanced data processing methods and powerful predictive models opens up a wide range of applications in the financial domain. By embracing these approaches, organizations can significantly improve their risk assessment capabilities and make data-driven decisions to mitigate potential losses and optimize their operations.